RC meeting and Run Plan, Wednesday Sept. 6, 2023

From SBS wiki
Jump to: navigation, search

Return to the RC Meeting Calendar

RC Meeting Information

Time:at 3:30pm today

  • Location: Counting house conference room
  • ZoomGov Meeting Connection Information


Useful Links

Readiness for Beam

Accelerator readiness:

  • CG: beam restoration to hall A now projected to start swing Tue Sept 12. We will need to have a shift crew starting on Wed Sept 13 day. Will Tireman to cancelling shifts.
  • SSG still has a number of items for the HCO but they will probably get to them by the end of the week (after completing Hall C work)

Hall readiness and lockup schedule (Zak Remele)

Beamline readiness: beamline quads and harps, bcms, bpms, ion chambers (Ciprian Gal)

  • Daniel Moser: Beam plans for reviving the target protection systems (the position locks and ramp rate) in place (Bill will review).
    • Beam current limit: RSAD written for 50uA.

Subsystems Checklist

For each subsystem we ask that the responsible person go over the following points:

  1. The overall status of your subsystem
  2. Will your subsystem be ready well before Hall lock-up on Friday or at the latest by Monday morning?
  3. Does your system being ready depend on your team only or are outside resources required?


  • HCal (Jiwan Poudel)
    • few channels with problems (probably cabling) and expect to recover them ahead of closeout on Friday
    • NEED update
  • Bigbite (Arun Tadepalli)
    • already took some cosmic data and things look ok
    • NEED update
  • Target cell status(Arun/Gordon)
    • Will wait for magnet upadte
  • Target laser status (alignment, polarization, power etc)
    • work ongoing
  • Counting House Alarms (Bill Henry)
    • Bill will test the alarms and instructions for shift crew response in the wiki
  • DAQ (Alex Camsonne)
    • SBS beamline crate recovery ongoing but expect to have it done by Friday
    • there seems to be a sitewide helicity signal issue; will talk to Paul King
  • GEMs (Holly Szumila-Vance)
    • Sean: SBS GEMs require debugging (but they should not hold off closing the hall)
  • Target magnets and measurements (Bogdan W. and Bill Henry)
    • Since SBS magnet is not working (temp fault) these tests could not be done.
  • Moller polarimeter (Don Jones)
    • Target motion not working (Ellen looking into it)
    • we want to do at least one measurement at the beginning; all systems that could be checked now (besides the target) work as expected
    • magnet not cold; expect to get it there end of week (over the weekend)
  • SBS magnet and Bigbite magnets ramp and hold?
    • ongoing issue (see above)
  • Corrector magnets:
    • need to be tested by Ellen; CG provided max current
  • Hodoscopes:
    • CG will check the elog

Plans and Procedures

Are required plans and procedures in place?

  • Is there a clear and up to date procedure for bring beam into the hall for the first time?(Eg. [1])
  • Ion chamber calibrations: Gen/GMn (PolHe3) in Hall A Fall 2023 Ion Chamber Calibrations -- https://tasklists.jlab.org/bslist/tasks/109834
  • Are the previous target protections still in place? Do we need to remind MCC to re-establish them?
    • Reminder: software GUIs with target position protection using BPMs engaged and turns beam off if it wanders +/1mm from set position. Must go to carbon target or target out to tune if beam trips due to this.
    • Hall A current ramp protection has maximum ramp rate and does what??? if it goes over a threshold above the set ramp rate.
    • Hall A Current Ramp Rate Protection Checkout and Commissioning -- https://tasklists.jlab.org/bslist/tasks/109836
    • Hall A BPM Target Protection Checkout -- https://tasklists.jlab.org/bslist/tasks/109835
  • Are the existing written procedures for re-establishing beam after being down for a period of time still up to date. Are the thresholds triggering the re-establishment procedure still valid?
  • Can we continue using the same spreadsheet?
  • Is there a clear experimental plan for the RC to accomplish for the first week of running? (Andrew Puckett and Bogdan Wojtsekhowski)
    • When will we do polarimetry? => early in the run to compare with hall C results
    • Are there optics runs we need to take?
    • What calibrations need to be done? Any sieve in data? -> Possibly.